OpenStudy (anonymous):

plz check my answer Solve 2^x + 3 = 7 x ≈ –5.19 x ≈ 0.94 ---> My Answer x ≈ –0.19

OpenStudy (anonymous):

No, it is wrong.

OpenStudy (anonymous):

\[2^{x}=4 => x \ln 2=\ln 4 => x = \frac{ \ln4 }{ \ln2 } =>\] solve you will get 2

OpenStudy (anonymous):

ok thanks